Two nearby trumpets are sounded together and a beat frequency of 8 Hz is heard. If one of the trumpets sounds at a frequency of 871 Hz, what are the two possible frequencies of the other trumpet? (Enter your answers from the smallest to the largest, in Hz, to three significant figures
Answer #1

Solution:

Beat frequency is defined as the absolute difference between the frequencies produced by the alternating constructive and destructive interference. It is expressed as:

fbeat = f1 - f2

Calculating the first possible frequency produced by the trumpet,

fbeat =| f1 - f2 |

or, 8Hz = 871 - f2

or, f2 = 863 Hz

Calculating the second possible frequency,

fbeat =| f1 - f2 |

Or, 8 Hz = f2 - 871

Or, f2 = 879 Hz

Hence, two possible frequencies are 863 Hz , and 879 Hz